Question 2. A doctor prescribes a sick freshman fruits (Banana and Orange) that provide Vitamin A, Vitamin B and Vitamin C. A pound of Banana contributes 8 grams of Vitamin A and 6 grams of Vitamin B and 2 grams of Vitamin C while a pound of Orange contributes 2 grams of Vitamin A and 6 grams of Vitamin B. A pound of Banana costs $4 and a pound of Orange costs $6. The student has to meet a minimum requirement of 27 grams of Vitamin A, 38 grams of Vitamin B and 6 grams of Vitamin C. The student should avoid taking more than 53 grams of Vitamin A and 71 grams of Vitamin B to avoid side effects such as nausea and vomiting. The pounds of banana eaten should be more than the pounds of orange eaten by at least 2 pound. Formulate a linear programming problem to help the poor freshman to choose the fruit mix to recover fast
Answer #1

Formulation of Linear Programming Problem:

  • Vitamin A
    (gm/pound)
  • Vitamin B
    (gm/pound)
  • Vitamin c
    (gm/pound)
  • Cost
    ($/pound)
  • Banana
    (X1)
8 6 2 4
  • Orange
    (X2)
2 6 0 6
  • Minimum
    Requirement
27 38 6
  • Maximum
    Requirement
53 71 -

So, we need to minimize the total cost of purchasing:

MINIMIZE Z = 4 X1 + 6 X2

Subject to Constraints:

  • 8 X1+ 2 X2 >= 27
  • 6 X1+ 6 X2 >= 38
  • 2 X1+ 0 X2 >= 6
  • 8 X1+ 2 X2 <= 53
  • 6 X1+ 6 X2 <= 71
  • X1 > 0 , X2> 0